The difference between two numbers is 9. The smaller number plus the larger is 25. What are the numbers?
Answer by blwinbbbles(106)

s = smaller number l = larger number
so the difference between the numbers is 9..so you can say:
l = s + 9
and the two added together is 25 ...so you can say:
l + s = 25..
since the first equation tells us what l is equal to (s + 9) we can substitute it in to the second equation
s + 9 + s = 25
2s + 9 = 25 substract 9 from both sides
2s = 16 divide both sides by 2
s = 8
Now that you know s you can plug it into the other equation..so:
l = 8 + 9
l = 17
So now you know that the smaller number is 8 and the larger number is 17